# Gauge preservation in renormalization for Yang-Mills and gravity theories

### v2

@@ 3.7 Physical interpretation of the counter terms for gravitational theories
$-\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})\widetilde{s}P_{f}^{\mu\nu}-L_{0\alpha}\widetilde{s}Q^{\alpha}_{f\beta}\eta^{\beta}]+O(\hbar^{2}).$ (3.144)
This action is written purely in terms of the first order redefined fields. In addition, it is BRST invariant at first order, since, as discussed above, the addition of the last two terms does not spoil BRST invariance due to the property $\widetilde{s}^{2}=0$. Furthermore $S_{(0)gf}$ is BRST invariant where now the BRST action affects only the finite part of the fields. The last effective action solves the Slavnov-Taylor and ghost identities by construction, up to violating terms that are of order 2superscriptPlanck-constant-over-$\hbar^{2}$ or higher. This is due to the fact it is designed as a solution of (3.128), that is, the Slavnov-Taylor equation at first order.
@@ 4. Discussion
In the present text, several important features about multiplicative renormalization in Yang-Mills theories and the renormalization of quadratic gravity were discussed, exploiting the analogies as much as possible. A key point for interpreting the counter terms appearing in the gravity case [6] is that the effective action order by order is expressed as the original classical action in terms of redefined fields, together with non standard though BRST invariant coupling to the ghost currents. These couplings are absent in Yang-Mills theories but they do not spoil BRST symmetry at a given order.
The discussion given along the text about renormalizabilty of quadratic gravity of course, does not cure the problem of unitarity of these theories. A possible and interesting investigation is to study how this procedure may be generalized to the models described in [8]-[21]. In particular the formulas (4.35)-(4.45) of reference [22] presents a path integral method for dealing with theories such as the present, which are of higher order and contain apparent instabilities. This method involves an additional factor in the measure of the path integral, indicated by $C$ in that reference, which involves the metric determinant and the volume of the space slices. It may be interesting to study how this factor modify the already existing renormalization formalism and if the renormalization program with this quantization procedure may be achieved. In any case, the generalization of these methods to that context may be of special interest.

### v3

@@ 3.7 Physical interpretation of the counter terms for gravitational theories
$-\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})\widetilde{s}P_{f}^{\rho\sigma}-L_{0\alpha}\widetilde{s}Q^{\alpha}_{f\beta}\eta^{\beta}]+O(\hbar^{2}).$ (3.146)
This action is written purely in terms of the first order redefined fields. In addition, it is BRST invariant at first order, since, as discussed above, the addition of the last two terms does not spoil BRST invariance due to the property $\widetilde{s}^{2}=0$. Furthermore $S_{(0)gf}$ is BRST invariant and the last effective action solves the Slavnov-Taylor and ghost identities by construction, up to violating terms that are of order 2superscriptPlanck-constant-over-$\hbar^{2}$ or higher. This is due to the fact it is designed as a solution of (3.130), that is, the Slavnov-Taylor equation at first order.
@@ 4. Discussion
In the present text, several important features about multiplicative renormalization in Yang-Mills theories and the renormalization of quadratic gravity were discussed, exploiting the analogies between both scenarios as much as possible. A key point for interpreting the counter terms appearing in the gravity case [6] is that the effective action is expressed, order by order in $\hbar$, as the original classical action in terms of redefined fields, together with non standard though BRST invariant coupling to the ghost currents. These couplings are absent in Yang-Mills theories, nevertheless they do not spoil BRST symmetry at a given order.
The discussion given along the text about renormalizabilty of quadratic gravity of course, does not cure the problem of unitarity of these theories. A possible and interesting investigation is to study how this procedure may be generalized to the models described in [8]-[21]. In particular the formulas (4.35)-(4.45) of reference [22] present a path integral method for dealing with theories such as the present, which are of higher order and contain apparent instabilities. This method involves an additional factor in the measure of the path integral, indicated by $C$ in that reference, which involves the metric determinant and the volume of the space slices. This factor does not affect perturbative calculation around flat spaces, as shown in [95], since they involve a term proportional to $\delta^{4}(0)$ which is cured by dimensional regularization. For perturbation around curved spaces, this determinant may play a role.